Given the data 6, 9, 12, 15, 16, 19, 22. Construct a sampling distribution of the sample mean without replacement and repetition by selecting 2 samples at a time.


1
Expert's answer
2022-02-22T11:56:48-0500

We have population values "6,9,12,15,16,19,22" population size "N=7"  and sample size "n=2."

Thus, the number of possible samples which can be drawn without replacements is "\dbinom{7}{2}=21."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ values & Sample\ mean\ (\bar{x}) \\ \hline
 6,9 & 7.5 \\
 \hdashline
 6,12 & 9 \\
 \hdashline
 6,15 & 10.5 \\
 \hdashline
 6,16 & 11 \\
 \hdashline
 6,19 & 12.5 \\
 \hdashline
 6,22 & 14 \\
 \hdashline
 9,12 & 10.5 \\
 \hdashline
 9,15 & 12 \\
 \hdashline
 9,16 & 12.5 \\
 \hdashline
 9,19 & 14 \\
 \hdashline
 9,22 & 15.5 \\
 \hdashline
 12,15 & 13.5 \\
 \hdashline
 12,16 & 14 \\
 \hdashline
 12,19 & 15.5 \\
 \hdashline
 12,22 & 17 \\
 \hdashline
 15,16 & 15.5 \\
 \hdashline
 15,19 & 17 \\
 \hdashline
 15,22 & 18.5 \\
 \hdashline
 16,19 & 17.5 \\
 \hdashline
 16,22 & 19 \\
 \hdashline
 19,22 & 20.5 \\
 \hdashline
\end{array}"



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 & \bar{x} & f & f(\bar{x}) & \bar{x}f(\bar{x})\\ \hline
 & 7.5 & 1 & 1/21 & 5/14 \\
 \hdashline
 & 9 & 1 & 1/21 & 3/7 \\
 \hdashline
 & 10.5 & 2 & 2/21 & 1 \\
 \hdashline
 & 11 & 1 & 1/21 & 11/21 \\
 \hdashline
 & 12 & 1 & 1/21 & 4/7 \\
 \hdashline
 & 12.5 & 2 & 2/21 & 25/21 \\
 \hdashline
 & 13.5 & 1 & 1/21 & 9/14 \\
 \hdashline
 & 14 & 3 & 1/7 & 2 \\
 \hdashline
 & 15.5 & 3 & 1/7 & 31/14 \\
 \hdashline
 & 17 & 2 & 2/21 & 34/21 \\
 \hdashline
 & 17.5 & 1 & 1/21 & 5/6 \\
 \hdashline
 & 18.5 & 1 & 1/21 & 37/42 \\
 \hdashline
 & 19 & 1 & 1/21 & 19/21 \\
 \hdashline
 & 20.5 & 1 & 1/21 & 41/42 \\
 \hdashline
 Sum= & & 21 & 1 & 99/7 \\
 \hdashline
\end{array}"


"\mu_{\bar{x}}=E(\bar{X})=99/7"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
x & p(x)\\ \hline
7.5 & 1/21\\
 \hdashline
9 & 1/21 \\
 \hdashline
10.5 & 2/21 \\
 \hdashline
11 & 1/21 \\
 \hdashline
12 & 1/21 \\
 \hdashline
12.5 & 2/21 \\
 \hdashline
 13.5 & 1/21\\
 \hdashline
 14 & 1/7\\
 \hdashline
 15.5 & 1/7 \\
 \hdashline
17& 2/21 \\
 \hdashline
17.5 & 1/21 \\
 \hdashline
18.5& 1/21 \\
 \hdashline
19& 1/21 \\
 \hdashline
20.5& 1/21 \\
\end{array}"